A biased coin is tossed. The probability of getting a head is 0.8 and the probability of getting a tail is 0.2. What is the probability of getting the 3rd head in the 13th throw?
For this problem, is it correct to make use of the Binomial distribution with n = 13 and r = 3?

Answer by stanbon(75887)

A biased coin is tossed. The probability of getting a head is 0.8 and the probability of getting a tail is 0.2. What is the probability of getting the 3rd head in the 13th throw?
--------
The pattern has 13 coin tosses with 3rd head in 13 position.
Example of the pattern:: hhtttttttttth
-----------------------
The two heads can be placed in 12C2 = (12*11)/(1*2) = 66 different ways
-----
Probability of each pattern is (1/2)^13
----
Ans: 66(1/2)^13 = 0.0081
------------------------------
Cheers,
Stan H.
==============================